I'm worse at geometry than I am at algebra, can someone please help me?
A rectangular pen for a pet is 3 feet longer than it is wide. Give possible values for the width, W of the pen if its area must be beyween 238 & 460 square feet inclusively.
I'd be grateful.

Answer by jorel1380(3719)

W(W+3)=238 to 460
W2+3W-238=0
(W+17)(W-14)=0
W=-17 or 14
W2+3W-460=0
(W+23)(W-20)=0
W=-23 or 20
--------------------
Throwing out the negative answers, we get the range of the width to be from 14 to 20 feet, inclusive..